OpenStudy (anonymous):

Solve A=h/2(b1 + b2) for h. Is this like the one before? We are a bit confused.

OpenStudy (turingtest):

this is\[A=\frac{h}{2}(b_1+b_2)\]right?

OpenStudy (anonymous):

Yes, we didn't know how to type it out like on the paper.

jhonyy9 (jhonyy9):

h=2A/(b1 +b2)

OpenStudy (turingtest):

1)multiply both sides by 2 2) divide both sides by (b1+b2) don't worry about the (b1+b2) it acts like a regular number x, as you can see in jhonyy's answer

OpenStudy (anonymous):

Okay, we'll try to write this out and show you how we did it just to double check.

OpenStudy (anonymous):

2A = h (b1+b2) <= Times by 2 2A = h (b1+b2) /(b1+b2) = 2A/(b1=b2) = h ?

OpenStudy (turingtest):

you have an equals sign where there should be a plus sign, but other than that, looks good

OpenStudy (anonymous):

So, there should be a plus sign between the 2A and the h?

OpenStudy (turingtest):

1) multiply both sides by 2\[2(A)=2\frac{h}{2}(b_1+b_2)\to 2A=h(b_1+b_2)\]2) divide both sides by (b1+b2)\[\frac{2A}{(b_1+b_2)}=\frac{h(b_1+b_2)}{(b_1+b_2)}=h\]

OpenStudy (anonymous):

Oh, okay. That looks a lot clearer. Thank you.

OpenStudy (turingtest):

np